# Ideal transition systems

### v2

@@ 6.2 Transition forests
It is not guaranteed that this process will yield a family satisfying every part of Definition 3.1: problems can occur with conditions (T3) and (T4). However, the construction of $\mathcal{T}_{\mathscr{I}}$ often does produce a transition system.
We visualize $\mathcal{T}_{\mathscr{I}}$ by drawing a forest graph whose vertices are the ideals $I$ with $(I,J)\in\mathcal{T}_{\mathscr{I}}$. In this transition forest, the children of a vertex labeled $I$ are the elements of $\mathscr{P}_{I}$. Although we have formulated these definitions specifically for the case $A=\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$, the relevant ideas easily extend to the case when $A$ is any affine variety.
@@
We revisit the ideals =$\mathscr{I}=\left\{I(\mathsf{MSV}^{\mathrm{sym}}_{w}):w\in I_{\infty}^{(2)}\right\}$ of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{2\times 2}]$ from Example 3.9. If we define $I_{1},\dots,I_{6}$ as in that example, then =$\mathscr{I}=\{I_{i}:i\in[5]\}$ and the transition forest of $\mathcal{T}_{\mathscr{I}}$ is
$I_{1}=\left\langle{u}_{21}^{2}-{u}_{11}{u}_{22}\right\rangle$$\boxed{I_{6}=\left\langle{u}_{11},\ {u}_{21}^{2}\right\rangle}$$I_{4}=\left\langle{u}_{21},\ {u}_{11}\right\rangle$$I_{5}=\left\langle{u}_{22},\ {u}_{21},\ {u}_{11}\right\rangle$   $I_{2}=0$$I_{3}=\left\langle{u}_{11}\right\rangle$$I_{4}=\left\langle{u}_{21},\ {u}_{11}\right\rangle$$I_{5}=\left\langle{u}_{22},\ {u}_{21},\ {u}_{11}\right\rangle$
@@
Consider the ideals =$\mathscr{I}=\left\{I(\mathsf{MSV}^{\mathrm{sym}}_{w}):w\in I_{\infty}^{(3)}\right\}$ of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{3\times 3}]$. In this case $\mathscr{I}$ has 14 elements $I_{1},I_{2},\dots,I_{14}$, which are indicated in Figure 1. We have used Macaulay2 to find one valid construction of $\mathcal{T}_{\mathscr{I}}$. This requires adding 7 more ideals $I_{15},I_{16},\dots,I_{21}$, also shown in Figure 1. Figure 2 displays the associated transition forest.
@@
$w\in I_{\infty}^{(3)}$ $I(\mathsf{MSV}^{\mathrm{sym}}_{w})\subset\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{3\times 3}]$
@@
[…]
@@
The family of ideals =$\mathscr{I}=\left\{I(\mathsf{MSV}^{\mathrm{sym}}_{w}):w\in I_{\infty}^{(4)}\right\}$ in $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{4\times 4}]$ has 43 elements. We have used Macaulay2 to construct $\mathcal{T}_{\mathscr{I}}$ and check that it is a transition system, verifying Conjecture 6.3 when $n=4$. Our calculation gives a set $\mathcal{T}_{\mathscr{I}}$ with 86 elements. The corresponding transition forest is too large to show here, but we mention that it has vertices with $\geq 2$ children.

### v3

@@ 6.2 Transition forests
It is not guaranteed that this process will yield a family satisfying every part of Definition 3.1: problems can occur with conditions (T3) and (T4). However, the construction of $\mathcal{T}_{\mathscr{I}}$ sometimes does produce a transition system.
We visualize $\mathcal{T}_{\mathscr{I}}$ by drawing a forest graph whose vertices are the ideals $I$ with $(I,J)\in\mathcal{T}_{\mathscr{I}}$. In this transition forest, the children of a vertex labeled $I$ are the elements of $\mathscr{P}_{I}$. Although we have formulated these definitions specifically for the case $A=\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$, the relevant ideas easily extend to the case when $A$ is any affine space.
@@
We revisit the ideals $\mathscr{I}=\left\{I(\mathsf{MSV}^{\mathrm{sym}}_{w}):w\in\mathcal{I}_{\infty}^{(2)}\right\}$ of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{2\times 2}]$ from Example 3.9. If we define $I_{1},\dots,I_{6}$ as in that example, then $\mathscr{I}=\{I_{i}:i\in[5]\}$ and the transition forest of $\mathcal{T}_{\mathscr{I}}$ is
[…]
@@
Consider the ideals $\mathscr{I}=\left\{I(\mathsf{MSV}^{\mathrm{sym}}_{w}):w\in\mathcal{I}_{\infty}^{(3)}\right\}$ of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{3\times 3}]$. In this case $\mathscr{I}$ has 14 elements $I_{1},I_{2},\dots,I_{14}$, which are indicated in Figure 1. We have used Macaulay2 to find one valid construction of $\mathcal{T}_{\mathscr{I}}$. This requires adding 7 more ideals $I_{15},I_{16},\dots,I_{21}$, also shown in Figure 1. Figure 2 displays the associated transition forest.
@@
$w\in\mathcal{I}_{\infty}^{(3)}$ $I(\mathsf{MSV}^{\mathrm{sym}}_{w})\subset\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{3\times 3}]$
@@
$I_{1}$$I_{6}$$I_{8}$$I_{10}$$I_{12}$$I_{13}$$I_{14}$$I_{2}$$I_{7}$$I_{9}$$I_{12}$$I_{13}$$I_{14}$$\boxed{I_{15}}$$I_{10}$$I_{12}$$I_{13}$$I_{14}$$I_{3}$$\boxed{I_{16}}$$I_{8}$$I_{10}$$I_{12}$$I_{13}$$I_{14}$$I_{4}$$I_{9}$$I_{12}$$I_{13}$$I_{14}$$\boxed{I_{17}}$$I_{10}$$I_{12}$$I_{13}$$I_{14}$$\boxed{I_{18}}$$I_{12}$$I_{13}$$I_{14}$$I_{5}$$\boxed{I_{19}}$$I_{11}$$\boxed{I_{20}}$$I_{13}$$I_{14}$$\boxed{I_{21}}$$\boxed{I_{20}}$$I_{13}$$I_{14}$
@@
The family of ideals $\mathscr{I}=\left\{I(\mathsf{MSV}^{\mathrm{sym}}_{w}):w\in\mathcal{I}_{\infty}^{(4)}\right\}$ in $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{4\times 4}]$ has 43 elements. We have used Macaulay2 to construct $\mathcal{T}_{\mathscr{I}}$ and check that it is a transition system, verifying Conjecture 6.3 when $n=4$. Our calculation gives a set $\mathcal{T}_{\mathscr{I}}$ with 86 elements. The corresponding transition forest is too large to show here, but we mention that it has vertices with $\geq 2$ children.